ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 121, 2001

Abramovich, D. and K. Matsuki Uniformity of stably integral points on principally polarized abelian varieties of dimension ≤ 2

351

Adams, S. and G. Stuck Isometric actions of SL_{n}(ℝ)×ℝ^{n} on Lorentz manifolds

93

Aljadeff, E. and D. Haile Division algebras with a projective basis 173

Bastardas, G. and C. Casacuberta A homotopy idempotent construction by means of simplicial groups

333

Casacuberta, C. See G. Bastardas

Chong, C. T., L. Qian, T. A. Slaman and Y. Yang Σ_{2} induction and infinite injury priority arguments, Part III: Prompt sets, minimal pairs and Shoenfield’s Conjecture

1

Danilenko, A. I. Funny rank-one weak mixing for nonsingular Abelian actions

29

de Jong, A. J. A conjecture on arithmetic fundamental groups  61

Dufour, J. P. and M. Zhitomirskii Singularities and bifurcations of 3-dimensional Poisson structures

199

Duval, A. M. and P. Zhang Iterated homology and decompositions of simplicial complexes

313

Grzeszczuk, P., A. Leroy and J. Matczuk Artinian property of constants of algebraic q-skew derivations

265

Gurevich, N. The relative trace formula for groups with involution  125

Haile, D. See E. Aljadeff

Kojman, M. Cantor–Bendixson degrees and convexity in ℝ^{2}  85

Leroy, A. See P. Grzeszczuk

Leung, D. H. The normed and Banach envelopes of weak L^{1}  247

Makar-Limanov, L. On the group of automorphisms of a surface

x^{n}y = P (z) 113

Matczuk, J. See P. Grzeszczuk

Matsuki, K. See D. Abramovich

McCutcheon, R. An infinitary polynomial Hales–Jewett theorem  157

Parent, P.-E. Homotopy localization functor L_{f} with respect to maps f having a wedge of spheres as homotopy cofibre 143 Qian, L. See C. T. Chong Rosenberg, D. and S. Sorin An operator approach to zero-sum

repeated games

221

Sarig, O. M. Thermodynamic formalism for null recurrent potentials 285

Slaman, T. A. See C. T. Chong Sorin, S. See D. Rosenberg Stuck, G. See S. Adams

Wang, Y. Bounded domains of the Fatou set of an entire function  55

Yang, Y. See C. T. Chong Zhang, P. See A. M. Duval Zhitomirskii, M. See J. P. Dufour